OpenStudy (anonymous):

Find the first quadrant area bounded by y=2^x and x=3.

OpenStudy (anonymous):

If this is a calculus problem, then you need to integrate the y=2^x function from x=0 to x=3

OpenStudy (anonymous):

yes it is thank you

OpenStudy (anonymous):

In case you are still confused...say 2^x=t and x=(1/ln2)*lnt. Then, the integral should be pretty straightforward from there.